OpenStudy (anonymous):

convert from regular form to polar form x^2 + xy + y^2 = 1

OpenStudy (anonymous):

Use the conversions: \[x=r\cos\theta\\ y=r\sin\theta\]

OpenStudy (anonymous):

I did and when I do I end up getting r^2 + r(cos(theta)sin(theta)) = 1 and I don't know where to go from here or if that's even right

OpenStudy (anonymous):

\[x^2+xy+y^2=1\\ (r\cos\theta)^2+(r\cos\theta)(r\sin\theta)+(r\sin\theta)^2=1\\ r^2\cos^2\theta+r^2\cos\theta\sin\theta+r^2\sin^2\theta=1\\ r^2\left(\cos^2\theta+\sin^2\theta +\cos\theta\sin\theta\right)=1\] Do you see what you can do with this?

OpenStudy (anonymous):

wouldn't you combine cos^2(theta) + sin^2(theta) to = 1

OpenStudy (anonymous):

Yep, that's right. So you have \[r^2\left(1+\cos\theta\sin\theta\right)=1\]

OpenStudy (anonymous):

so what would be the next step?

OpenStudy (anonymous):

Solve for r. Or rather, r².